OpenStudy (anonymous):

Two tug boats are towing a large boat, of mass M kg, back to shore. Tug boat 1 is pulling with a force of T1 N at an angle of 25 degrees north of the forward motion and tug boat 2 is pulling with a force of T2 N at an angle of 25 degrees south of the forward motion. If the large boat is being pulled with constant velocity, and there is a resistive force of 4000N to the motion, then what are the magnitudes of the two forces T1 and T2?

OpenStudy (mathmate):

Large boat has constant velocity, so only resistive force is F=4000N It will be assumed that the large boat has a net velocity due east (since tug boats are symmetrically placed north and south). Then sum forces N-S, or Fy: Sum Fy=0 => T1sin(theta1)+T2sin(theta2)=0 => T1=T2 Sum Fx=0 => T1cos(theta1)+T2cos(theta2)=F Solve for T1 and T2.

OpenStudy (anonymous):

Please show steps! Im confused

OpenStudy (mathmate):

Can you start by drawing a diagram, this is the first step to solving the problem.

OpenStudy (anonymous):

Im on my phone right now, so I cant

OpenStudy (mathmate):

That's OK, take your time. The best way to solve a physics (statics) problem is to visualize the problem in relation to its parameters. This is most often done by a diagram indicating the object in question, forces acting on it, the direction, magnitude, and the notation you use. Finally, you'd draw a free-body diagram and calculate forces by summing in each direction.